OpenStudy (anonymous):

HELP! THIS THE THIRD TIME I HAVE BUMBED THIS. Help appreciated! What is the absolute min/max of x(sqrt(x^2+16)) I got this derivative sqrt(x^2+16)+(x^2)/(sqrt(x+16), which you set to zero to find the min/max. I cant seem to find them!

OpenStudy (neer2890):

is this your function: \[f(x)=x \sqrt{x ^{2}+16}\]

OpenStudy (anonymous):

yes!

OpenStudy (neer2890):

\[f \prime(x)=\sqrt{x ^{2}+16}+\frac{ x }{ 2\sqrt{x ^{2}+16} }\]=0

OpenStudy (neer2890):

take l.c.m \[\frac{ 2(x ^{2}+16)+x }{ 2\sqrt{x ^{2}+16} }=0\]

OpenStudy (anonymous):

I have and I cant seem to find the min/ or max I keep getting + -2sqrt2

OpenStudy (neer2890):

\[2x^{2}+x+32=0\] solve for x

OpenStudy (neer2890):

okay. you want me to solve this further.

OpenStudy (anonymous):

yes please!

OpenStudy (neer2890):

is there any range in which this function is defined?

OpenStudy (anonymous):

(-5,7)

OpenStudy (neer2890):

can you tell me how you got \[\pm2\sqrt{2}\]

OpenStudy (anonymous):

honestly, it was wolframalpha since, I cant seem to get through my arithmetic fully

OpenStudy (neer2890):

okay, just find out the roots of that equation and plug in all four values (2 values you will find out and 2 values are -5 and 7) in \[f \prime(x)\] and check on which value the function is minimum and maximum.

OpenStudy (neer2890):

the value at which the function is minimum is absolute minima and the value at which the function is maximum is absolute maxima. hope this will help.

OpenStudy (anonymous):

yes, this is what im asking, I cant seem to find the roots

OpenStudy (neer2890):

sorry., you have to plug in all 4 values in f(x) not in \[f \prime(x)\]

OpenStudy (neer2890):

okay. use this formula \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (neer2890):

here the roots of your equation are imaginary.So you have to find the critical points of your function where it becomes zero.

OpenStudy (anonymous):

ok so do I just se the equation to zero? to f prime?

OpenStudy (anonymous):

sorry I cant seem to find it thank u tho.

OpenStudy (neer2890):

can you help here... @hartnn

OpenStudy (triciaal):

@neer2890 i think you made an error when you found the derrivative. you forgot to "do the inside" the x^2 which would be 2x (x^2 + 16)^1/2 =1/2(x^2 +16)^-1/2*2x = x / (x^2 + 16)^1/2

OpenStudy (neer2890):

oh... thanks for correcting me @triciaal

OpenStudy (neer2890):

but the question still remains same. how to find the critical roots?

hartnn (hartnn):

method : equate the derivative to 0 if you don't find any real roots, then use the end-points plug in x = -5 and x =7 the smaller value will be miminum larger value will be maximum

OpenStudy (neer2890):

\[f \prime(x)=\frac{ 2x ^{2}+16 }{ \sqrt{x ^{2}+16} }=0\]

OpenStudy (neer2890):

so the absolute minima=-32 and absolute maxima=56.4 is it?

hartnn (hartnn):

its like this function,  minimum and maximum are just endpoints